Circular No. 5976

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994L IN NGC 2848 A. V. Filippenko and L. C. Ho, University of California at Berkeley,

report that CCD spectra (range 310–800 nm, resolution 0.6 nm), obtained on Apr. 18 UT with the Shane 3-m reflector at Lick Observatory, confirm that SN 1994L (cf. IAUC 5975) is of type II, with prominent P-Cyg profiles. The recession velocity of a superposed H II region is 2070 km/s. The supernova is ∼ 13′′ west and 56′′ south of the galaxy nucleus.

PECULIAR VARIABLE IN SAGITTARIUS M. L. Hazen, Harvard-Smithsonian Center for Astrophysics, reports

that an examination of 272 blue plates of the field of this object (IAUC 5961), distributed from 1888 to 1989 with a few gaps, shows no evidence of this variable to a limiting magnitude of B ∼ 14.5 (or, in some cases, considerably fainter).

SUPERNOVA 1994D IN NGC 4526 R. W. Argyle and L. V. Morrison, Royal Greenwich Observatory, report

an accurate optical position and photometry for SN 1994D obtained with the Carlsberg Automatic Meridian Circle (operated by T. Knudsen, O. Einicke and L. Helmer on La Palma): α = 12^{h}34^{m}02^{s}.395 ± 0^{s}.003, δ = +7^{o}42′05′′.07 ± 0′′.06 (equinox FK_{5}/J2000.0; from 18 observations, epoch 1994.24). V magnitudes (± 0.03 mag): Mar. 16.09 UT, 12.16; 17.09, 12.09; 19.08, 11.85; 21.08, 11.85; 23.07, 11.86; 25.06, 11.84; 26.06, 11.85; 29.05, 12.08; 30.05, 12.30; Apr. 6.03, 12.66; 8.03, 12.89; 11.02, 13.12.

B. and J. Nicolet, Observatoire de Genève, report photometry obtained at the European Southern Observatory (Geneva photometric system): “Apr. 1.10 UT, V = 11.769, [V − B] = +0.252, [U − B] = +1.567; 2.03, 11.723, +0.143, +1.561; 6.02, 11.947, −0.184, +1.772. The colors are comparable to those of SN 1987A four days after explosion for the Apr. 1 and 2 measurements; the last measurement gives colors more or less comparable to SN 1987A + 6 days.”

Further CCD photometry by H. Mikuž, Ljubljana, Slovenia (cf. IAUC 5958): Mar. 29.97 UT, V = 12.11 ± 0.02, B − V = +0.31 ± 0.04, V − R = 0.00 ± 0.03; 30.93, V = 12.21 ± 0.02, B − V = +0.31 ± 0.04, V − R = +0.01 ± 0.03, V − I = −0.06 ± 0.03; Apr. 7.89, V = 12.71 ± 0.02, B − V = +0.82 ± 0.05.

1994 April 18 Daniel W. E. Green